OpenStudy (anonymous):

Help please, anyone? Sammy has an annuity that pays him $9600 at the beginning of each year. Assume the economy will grow at a rate of 3.1% annually. What is the value of the annuity if he received it now instead of over a period of 10 years? $78,044.65 $80,651.16 $81,075.98 $83,999.29

OpenStudy (anonymous):

@Daveqwerty Can you help me?

OpenStudy (anonymous):

I wish I could hun....I guess ill try.

OpenStudy (shamil98):

Is this for Algebra 2?

OpenStudy (anonymous):

Algebra with financial applications

OpenStudy (shamil98):

Oh, I just checked Pert, isn't used here, my bad.

OpenStudy (anonymous):

@shamil98 so then how is it solved?

OpenStudy (mathstudent55):

This is an annuity due (that pays at the beginning of each period). \(P = A \times \dfrac{(1 + i)^n - 1}{i} \times (1 + i) \) \(P = $9,600 \times \dfrac{(1 + 0.031)^{10} - 1}{0.031} \times (1 + 0.031) \) \(P = $113,988.83\)

OpenStudy (anonymous):

@mathstudent55 so then how do I find the annuity over 10 years?

OpenStudy (tkhunny):

"economy will grow"?? What kind fo information is that? It's NOT an interest rate.

OpenStudy (anonymous):

@tkhunny :(

OpenStudy (tkhunny):

We may just be discounting for inflation, I suppose, but we should point out that the interest rate is 0% while we're at it. That may or may not be as intended.

OpenStudy (anonymous):

@tkhunny This is a question, it's not based on reality

OpenStudy (tkhunny):

It has to make sense. Reality isn't all that related to making sense, either.

OpenStudy (tkhunny):

10 years of 9600 ==> 10*9600 = 96000 Okay, with no discounting for anything, we get $96,000. Therefore, if we discount AT ALL, we should get less than that, agreed?

OpenStudy (anonymous):

@tkhunny Yes, that does make sense

OpenStudy (tkhunny):

Above, we see that mathstudent was accumulating, and not discounting. That's why the result was greater than 96,000. If we have i = 0.031 -- (inflation rate for one year) And we define v = 1/(1+i) -- Annual Discount factor related to inflation. We have the Present Value of our ten payments described like this: 9600(1 + v + v^2 + ... + v^9) Agreed?

OpenStudy (mathstudent55):

Good point, @tkhunny. I saw "grow at a rate of 3.1% annually", and I thought it meant the investment grows at that rate. How about using the formula above but with a negative rate of growth, -3.1% = -0.031. Would it work? \(P = $9,600 \times \dfrac{(1 - 0.031)^{10} - 1}{-0.031} \times (1 - 0.031)\) \(P = $81,063.82\)

OpenStudy (tkhunny):

No, not quite the same thing. 1 - 0.031 = 0.969 1/1.031 = 0.9699321 Definitely close, though. 9600(1 + v + v^2 + ... + v^9) = \(9600\dfrac{1-v^{10}}{1-v}\) Go!

OpenStudy (mathstudent55):

@tkhunny Thanks for the explanation. I had never learned the formula for the annual discount factor related to inflation.

OpenStudy (tkhunny):

It IS confusing as it is presented. With just inflation, and no interest, it just isn't quite the same thing.

OpenStudy (mathstudent55):

I agree.